# Diffusion-Aided Joint Source Channel Coding For High Realism Wireless Image Transmission

### v2

@@ III-C Performance on General Image Transmission
In Figure 7, we assess the DiffJSCC framework using the ADE20K dataset at transmission rates $R=1/128$ and $R=1/32$, examining FID scores and mIoU for semantic segmentation performance. For mIoU evaluations, semantic masks are inferred from the reconstructed images utilizing a pre-trained ViT-adaptor [48]. The data indicates that while the DeepJSCC-$adv$ approach can effectively reduce FID scores, it comes at the cost of decreased mIoU, thus limiting its practicality for subsequent image processing tasks. In contrast, DiffJSCC notably improves both FID and mIoU. When $R=1/128$, it demonstrates an average improvement (reduction) in FID by $69.2\%$ and a $52.5\%$ increase in mIoU over DeepJSCC, alongside the notable improvement of $75.6\%$ in FID and $161\%$ in mIoU compared to DeepJSCC-$adv$. For the higher transmission rate $R=1/32$, DiffJSCC consistently shows superior performance, with a mean enhancement of $42.9\%$ in FID and $8.4\%$ in mIoU over DeepJSCC, along with gains of $33.3\%$ in FID and $19.1\%$ in mIoU in comparison to DeepJSCC-$adv$. With increasing SNR values, the mIoU obtained by DiffJSCC begins to align closely with that of the original undistorted images. Figure 8 presents a visual example from the ADE20K dataset. The example shows that, with DeepJSCC, elements such as grass and buildings are somewhat recognizable, yet roads and lamps appear unclear. The reconstructions obtained from DeepJSCC-$adv$ significantly lack authenticity, and the ViT-adaptor fails to generate accurate segmentation masks. On the other hand, images produced by DiffJSCC demonstrate higher visual quality to produce more accurate segmentation labels in most regions, thus highlighting the framework’s applicability for practical JSCC scenarios.
@@ III-D Performance on Facial Image Transmission
Figure 9 presents a comparison of our DiffJSCC framework with DeepJSCC and other generative JSCC methods, including InverseJSCC [33] and the approach presented by Yilmaz et al. [35]. For DiffJSCC, we showcase results with varying $\lambda$ values of 0, 10, and 50. The outcomes of InverseJSCC and Yilmaz et al. are taken from their original papers. We only evaluate the PSNR and LPIPS metrics since MS-SSIM and FID evaluations were not conducted in their works. Overall, DiffJSCC shows superior performance, surpassing Inverse-JSCC regarding both PSNR and LPIPS metrics. In comparison with the Yilmaz et al. method, DiffJSCC has a slightly lower PSNR but achieves superior LPIPS scores. Unlike the Kodak image results, where $\lambda=0$ provided the optimal LPIPS, this configuration introduces artifacts because of the significant domain gap between the Stable Diffusion training dataset and the CelebAHQ dataset. However, by utilizing direct guidance with the initial JSCC reconstructions (with $\lambda=50$), these artifacts are effectively mitigated, resulting in enhanced LPIPS performance. Quantitatively, at $\lambda=50$, DiffJSCC accomplishes an average enhancement of $15.9\%$ and $28.6\%$ over InverseJSCC for $R=1/256$ and $R=1/64$, respectively, and showcases an improvement of $10.3\%$ and $9.4\%$ for $R=1/256$ and $R=1/64$ against the Yilmaz et al. method. Figure 10 visualizes examples from the CelebA dataset, highlighting DiffJSCC’s ability to reconstruct high-quality images in extremely challenging conditions such as a very low bit rate ($R=1/256$) and a low SNR value ($\text{SNR}=-5\text{dB}$). As the SNR increases, the quality of the reconstructions improves, making them more closely resemble the original images.
@@ III-E1 The effect of different conditions
In this analysis, we investigate the impact of various conditional signals on the control module’s performance, as shown in Figure 11 bottom, emphasizing the importance of including multimodal information during the image restoration process. We observe that relying exclusively on spatial features extracted from the initial JSCC reconstruction offers limited effectiveness, whereas augmenting the model with channel SNR information yields an average FID score improvement of $2\%$, a benefit that becomes more pronounced at reduced SNR levels. This enhancement underlines the critical role channel state data play in enabling the denoiser to adjust to varying levels of distortion. Significantly, the incorporation of captions as textual prompts into the denoising framework further elevates FID performance by an additional $4.3\%$. Figure 11 left showcases that despite a natural decline in caption quality, quantified by the CLIP score[43], as SNR and transmission rates decrease, the presence of pivotal image keywords ensures sustained improvements, even under constrained conditions. This outcome illustrates the resilience and utility of textual cues in the denoising process. The benefit is further demonstrated by the marked performance increase when using captions derived from undistorted images, as indicated by the orange line in Figure 11 right.
Figure 12 serves as an excellent example of the influence of textual cues: The BLIPv2 framework is employed to generate coherent text prompts that capture the central themes of the image, despite the original JSCC reconstruction being affected by blur and distortion. Without textual cues, our approach can improve the general image quality but struggles to accurately recreate the parrot on the left side of the image. Incorporating text prompts addresses this issue significantly, enabling the precise restoration of intricate details like the parrot’s head, thereby demonstrating the crucial importance of textual information in the image reconstruction process.
@@
Figure 13: Effect of DDPM sampling steps for various SNRs with $R=1/128$.
III-E2 The effect of sampling steps
Our research investigates the impact of the number of denoising sampling steps on LPIPS and FID scores using the Kodak dataset with a transmission rate of $R=1/128$. The evolution of performance across different SNR levels is illustrated in Figure 13. It is evident that both LPIPS and FID metrics enhance considerably with an increase in the number of sampling steps. Notably, after reaching 50 steps, these enhancements level off, signifying reduced gains with additional steps across all evaluated SNR values. Using the specific example of $\text{SNR}=13\text{dB}$, increasing the sampling from 25 to 50 steps results in substantial gains of $0.02$ for LPIPS and $6.5$ for FID, whereas increasing from 50 to 100 steps yields more modest improvements of $0.004$ for LPIPS and $2.5$ for FID. These findings strongly support our decision to use 50 sampling steps in the experiments. This number is optimal because it saves computational resources, shortens inference time, and still achieves satisfactory performance, effectively balancing efficiency and effectiveness in our method.
@@ IV Conclusion
In this study, we introduce DiffJSCC, a novel deep JSCC framework that capitalizes on the robust priors of the pre-trained Stable Diffusion model to substantially improve the perceptual quality of reconstructed images. Unlike the traditional end-to-end learning approach, DiffJSCC initially derives multimodal conditions from the received channel symbols. For enhanced training stability and efficiency, DiffJSCC employs a JSCC decoder to create an initial reconstruction of the transmitted image, which is then utilized to generate multimodal features. By leveraging multimodal conditions, DiffJSCC effectively integrates spatial, textual, and channel state information to steer the image generation process. A control module is developed to fine-tune the Stable Diffusion model based on the generated conditions. Moreover, the initial JSCC reconstruction can also be utilized to guide the denoising process, balancing fidelity and realism. Extensive experimentation on both general and facial image transmission demonstrates that DiffJSCC considerably outperforms current generative JSCC methods, providing superior human-perceptual image quality and downstream task performance, particularly in environments with low transmission rates and poor SNRs. Additionally, the proposed DiffJSCC is model-agnostic and can be adapted to more advanced deep JSCC architectures or improved conditional diffusion models. Furthermore, DiffJSCC introduces no extra computations on the transmitter side, which is particularly beneficial for energy-limited scenarios.
Building upon the mentioned topics, there are several prospective future directions. For instance, it is possible to investigate more intricate channel models (such as multipath fading channels) and integrate more sophisticated channel state information into the control module. Furthermore, as the multimodal conditions stem from an initial JSCC reconstruction, their quality is bound by the JSCC decoder’s performance. Exploring alternative feature extraction methods that do not have this limitation would be worthwhile. Lastly, examining methods to minimize denoising steps to enhance efficiency on the receiver’s end would also be intriguing.

### v3

@@ III-C Performance on General Image Transmission
In Figure 8, we assess the DiffJSCC framework using the ADE20K dataset at transmission rates $\rho=1/384$ and $\rho=1/96$ under the AWGN channel, examining FID scores and mIoU for semantic segmentation performance. For mIoU evaluations, semantic masks are inferred from the reconstructed images utilizing a pre-trained ViT-adaptor [51]. The data indicates that while the DeepJSCC-$adv$ approach can effectively reduce FID scores, it comes at the cost of decreased mIoU, thus limiting its practicality for subsequent image processing tasks. In contrast, DiffJSCC notably improves both FID and mIoU. When $\rho=1/384$, it demonstrates an average improvement (reduction) in FID by $69.2\%$ and a $52.5\%$ increase in mIoU over DeepJSCC, alongside the notable improvement of $75.6\%$ in FID and $161\%$ in mIoU compared to DeepJSCC-$adv$. For the higher transmission rate $\rho=1/96$, DiffJSCC consistently shows superior performance, with a mean enhancement of $42.9\%$ in FID and $8.4\%$ in mIoU over DeepJSCC, along with gains of $33.3\%$ in FID and $19.1\%$ in mIoU in comparison to DeepJSCC-$adv$. With increasing SNR values, the mIoU obtained by DiffJSCC begins to align closely with that of the original undistorted images. Figure 9 presents a visual example from the ADE20K dataset. The example shows that, with DeepJSCC, elements such as grass and buildings are somewhat recognizable, yet roads and lamps appear unclear. The reconstructions obtained from DeepJSCC-$adv$ significantly lack authenticity, and the ViT-adaptor fails to generate accurate segmentation masks. On the other hand, images produced by DiffJSCC demonstrate higher visual quality to produce more accurate segmentation labels in most regions, thus highlighting the framework’s applicability for practical JSCC scenarios.
@@ III-D Performance on Facial Image Transmission
Figure 10 presents a comparison of our DiffJSCC framework with DeepJSCC and other generative JSCC methods, including InverseJSCC [35] and the approach presented by Yilmaz et al. [37] on CelebAHQ dataset. For DiffJSCC, we showcase results with varying $\lambda$ values of 0, 10, and 50. The outcomes of InverseJSCC and Yilmaz et al. are taken from their original papers. We only evaluate the PSNR and LPIPS metrics since MS-SSIM and FID evaluations were not conducted in their works. Overall, DiffJSCC shows superior performance, surpassing Inverse-JSCC regarding both PSNR and LPIPS metrics. In comparison with the Yilmaz et al. method, DiffJSCC has a slightly lower PSNR but achieves superior LPIPS scores. Unlike the Kodak image results, where $\lambda=0$ provided the optimal LPIPS, this configuration introduces artifacts because of the significant domain gap between the Stable Diffusion training dataset and the CelebAHQ dataset. However, by utilizing direct guidance with the initial JSCC reconstructions (with $\lambda=50$), these artifacts are effectively mitigated, resulting in enhanced LPIPS performance. Quantitatively, at $\lambda=50$, DiffJSCC accomplishes an average enhancement of $15.9\%$ and $28.6\%$ over InverseJSCC for $\rho=1/768$ and $\rho=1/192$, respectively, and showcases an improvement of $10.3\%$ and $9.4\%$ for $\rho=1/768$ and $\rho=1/192$ against the Yilmaz et al. method. Figure 11 visualizes examples from the CelebA dataset, highlighting DiffJSCC’s ability to reconstruct high-quality images in extremely challenging conditions such as a very low bit rate ($\rho=1/768$) and a low SNR value ($\text{SNR}=-5\text{dB}$). As the SNR increases, the quality of the reconstructions improves, making them more closely resemble the original images.
@@ III-E1 The effect of different conditions
In this analysis, we investigate the impact of various conditional signals on the control module’s performance on the Kodak dataset under the AWGN channel, as shown in Figure 12 right, emphasizing the importance of including multimodal information during the image restoration process. We observe that relying exclusively on spatial features extracted from the initial JSCC reconstruction offers limited effectiveness, whereas augmenting the model with channel SNR information yields an average FID score improvement of $2\%$, a benefit that becomes more pronounced at reduced SNR levels. Incorporating captions as textual prompts into the denoising framework further elevates FID performance by an additional $4.3\%$. Figure 12 left showcases that despite a natural decline in caption quality measured by the CLIP score[46] as SNR and transmission rates decrease, the presence of image keywords ensures sustained improvements in FID as shown in Figure 12 right. This outcome illustrates the resilience and utility of textual cues in the denoising process. The benefit is further demonstrated by the marked performance increase when using captions derived from undistorted images, as indicated by the orange line in Figure 12 right.
@@
Figure 14: Effect of DDPM sampling steps with $\rho=1/384$.
Figure 13 serves as a visualization example of the influence of textual cues: Despite the original JSCC reconstruction being affected by blur and distortion, The BLIPv2 model employed in our framework can generate coherent text prompts that capture the main context of the image. Without textual cues (Fig. 13 bottom left), our approach can improve the general image quality but struggles to accurately recreate the parrot on the left. Incorporating text prompts addresses this issue significantly (Fig. 13 bottom right), enabling the realistic restoration of intricate details like the parrot’s head, demonstrating the crucial importance of textual information in the image reconstruction process.
III-E2 The effect of the number of sampling steps
This section investigates the impact of the number of denoising sampling steps on LPIPS and FID scores using the Kodak dataset with a transmission rate of $\rho=1/384$. The performance evolution across different SNR levels is illustrated in Figure 14. It is evident that both LPIPS and FID metrics enhance considerably as the number of sampling steps increases. Notably, after reaching 50 steps, these enhancements level off with marginalized gains from additional steps across all evaluated SNR values. When $\text{SNR}=13\text{dB}$, for example, increasing the sampling from 25 to 50 steps results in substantial gains of $0.02$ for LPIPS and $6.5$ for FID, whereas increasing it from 50 to 100 steps yields reduced improvements of $0.004$ for LPIPS and $2.5$ for FID. These findings led to our decision to use 50 sampling steps in our experiments in earlier sections. Using 50 steps balances the efficiency and effectiveness of our method considering the required computational resources, inference time, and reconstruction image quality.
@@ IV Conclusion
In this study, we introduce DiffJSCC, a novel deep JSCC framework that capitalizes on the robust priors of the pre-trained Stable Diffusion model to substantially improve the perceptual quality of reconstructed images. Unlike the traditional end-to-end learning approach, DiffJSCC initially derives multimodal conditions from the received symbols. For enhanced training stability and efficiency, DiffJSCC employs a JSCC decoder to create an initial reconstruction of the transmitted image, which is then utilized to generate multimodal features. DiffJSCC effectively integrates multimodal conditions that consist of spatial, textual, and channel state information to steer the image generation process. Our framework employs a new control module to fine-tune the Stable Diffusion model based on the generated conditions, and it leverages the initial JSCC reconstruction to guide the denoising diffusion process, balancing fidelity and realism. Extensive experiments on both general and facial image transmission cases demonstrate that DiffJSCC considerably outperforms the latest generative JSCC methods, providing superior human-perceptual image quality and downstream task performance, particularly in environments with low transmission rates and poor SNRs. The proposed DiffJSCC is model-agnostic and can be extended with more advanced deep JSCC architectures and/or improved conditional diffusion models. Furthermore, DiffJSCC introduces no extra computations on the transmitter side, which is particularly beneficial for scenarios with energy-limited image-capturing devices.